• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name hyde

    Log entry time 23:49:36 on March13,2009

    Entry number 265993

    keyword=Eve 13 Mar 09 shift summary

    Shift Summary
    Evening Shift Friday the 13th of March 2009

    Exp E06-014
    Started at Kin 4E with 0.113 C accumulated in day shif
    Total goal = 0.49C

    Just started
    HRS-L run# 20662
    BB run# 2133
    spot++ raster size is 3x3 mm2

    17:00 End runs
    HRSL 20662 358K events, 0.046 C; and
    BB 2133 6.5M events, 0.046 C


    17:01 Start runs
    HRSL 20663
    BB 2164
    18:00 End runs
    HRSL 20663 350K, 0.046 C
    BB 2134 6.5M

    18:01 Start runs
    HRSL 20664
    BB 2135

    18;17 increase current from 14.2 muC to 15muC
    HRS-L DT = 3%
    BB DT = 16%

    18:29 End Runs
    HRSL 20664 0.021 C
    BB 2135

    Total Kin4E this shift 0.113

    18;30 3 minutes for Bogdan W rate test of GEM
    Junk monitor runs 20665 2166

    18:35 Start Runs (15 muA)
    HRSL 20666
    BB 2137
    19:25 end run 0.045 C

    19:26 Start Runs
    HRSL 20667
    BB 2138
    20:16 end runs 0.045 C

    20:17 Start Runs
    HRSL 20668
    BB 2139
    21:16 End Runs 0.045 C, Total at 4E = 0.355

    21:17 Start Runs
    HRSL 20669
    BB 2140
    21:48 End Runs

    Switch to Kin 4C
    HRS-L = 0.79998 GeV/c
    switch to N2 Reference Cell 5muA

    30 min of HRS-L N2 data run 29670
    BB coda crash, no N2 data

    22:40 back to 3He production kin 4C
    Start run
    HRS-L 20671
    BB 2141
    23:30 End runs 0.044 C

    23:31 Start new runs
    no beam
    MCC called, RF problem "few more minutes"
    Noted BCA flat lined for 25 minutes then spontaneously recovered 5 min
    before RF trip

    End of Shift
    Beam back at 15 muA
    BCA display OK
    run
    HRSL 20672
    BB 2142